Miller Company’s total sales are $204,000. The company’s direct labor cost is $24,480, which represents 30% of its total conversion cost and 40% of its total prime cost. Its total selling and administrative expense is $30,600 and its only variable selling and administrative expense is a sales commission of 5% of sales. The company maintains no beginning or ending inventories and its manufacturing overhead costs are entirely fixed costs.

Required:

1. What is the total manufacturing overhead cost?

2. What is the total direct materials cost?

3. What is the total manufacturing cost?

4. What is the total variable selling and administrative cost?

5. What is the total variable cost?

6. What is the total fixed cost?

7. What is the total contribution margin?
Answer #1
1
Total conversion cost 81600 =24480/30%
Total manufacturing overhead cost 57120 =81600-24480
2
Total prime cost 61200 =24480/40%
Total direct materials cost 36720 =61200-24480
3
Direct materials cost 36720
Direct labor cost 24480
Manufacturing overhead cost 57120
Total manufacturing cost 118320
4
Total variable selling and administrative cost 10200 =204000*5%
5
Direct materials cost 36720
Direct labor cost 24480
Variable selling and administrative cost 10200
Total variable cost 71400
6
Manufacturing overhead cost 57120
Fixed selling and administrative expense 20400 =30600-10200
Total fixed cost 77520
7
Sales revenue 204000
Less: Total variable cost 71400
Total contribution margin 132600